OpenStudy (anonymous):

How do you find three complex cubed roots of 6????

OpenStudy (anonymous):

Well a complex number is "i" so we're doing with that right?

OpenStudy (anonymous):

i believe so, i am just using this to study for my final and this was a question on the study guide that idk how to do

OpenStudy (phi):

you could write 6 in polar form \[ 6 = 6 e^{i 2 \pi} \] now raise to the 1/3 power: \[ 6^{\frac{1}{3} } e^{i \frac{2 \pi}{3} } \] to get the other roots notice you could start with \[ 6 = 6 e^{i 4 \pi} \] and do the same thing also \[ 6 = 6 e^{i 6 \pi} \] doing it again will start repeating the roots

OpenStudy (phi):

See http://tutorial.math.lamar.edu/Extras/ComplexPrimer/Roots.aspx for more background

OpenStudy (anonymous):

so would the roots be for example 6=6e^i6π or no? @phi

OpenStudy (phi):

to the 1/3 power

OpenStudy (phi):

the idea is to write 6 as \[ 6 e^{i 2 \pi k} \] where k is an integer we can use euler's formula to show this is \[ 6( \cos (2 \pi k) + i \sin(2 \pi k)) \] for k=0 this is \[ 6( \cos (0) + i \sin(0)) = 6(1+0) = 6 \] for k= 1 we still will get 6 (because 2 pi is the same angle as 0 ) same for k=2, and so on... now raise 6 to the 1/3 power. you get \[ 6^{\frac{1}{3}} e^{i \frac{2 \pi k}{3}} \] for k=0,1,2 (or even negative integer k's)

OpenStudy (phi):

you might think there are an infinite number of roots because k goes forever, but k=0 gives the same answer as k=3, and k=1 gives the same answer as k=4, and so on.... you get 3 distinct roots

OpenStudy (phi):

we start with 6 written as \[ \left( 6 e^{i 2 \pi k}\right) , \text{ k is an integer }\] and raise it the to 1/3 power \[ \left( 6 e^{i 2 \pi k}\right)^{\frac{1}{3}} \] which can be written as \[ 6^{\frac{1}{3}} e^{i \frac{2 \pi k}{3}} \]

OpenStudy (phi):

you can write the roots in rectangular form \[ 6^{\frac{1}{3}} \left(\cos(0) + i \sin(0)\right)\] \[ 6^{\frac{1}{3}} \left(\cos(120º) + i \sin(120º)\right)\] \[ 6^{\frac{1}{3}} \left(\cos(240º) + i \sin(240º)\right)\] notice we add 120 degrees to get to the next root. also, the next root 240+120= 360 gets us back to 0 degrees

OpenStudy (anonymous):

so for k=2 i would do 6(cos(2π(2))+isin(2π(2))) --> 6(cos(4π)+isin(4π)) --> 6((1) + i(-2*10^-13)) I feel like this is wrong.. @phi

OpenStudy (phi):

it looks like you used a calculator that 10^-13 is a tiny number, and should be exactly 0 you just found that 6 can be written as 6*( cos (4 pi) + i sin( 4 pi) ) see http://www.wolframalpha.com/input/?i=6*%28+cos+%284+pi%29+%2B+i+sin%28+4+pi%29+%29 to find the cube root (i.e. raised to the 1/3 power), you have to raise 6 to the 1/3 power \[ \left( 6 e^{i 2 \pi k}\right)^{\frac{1}{3}} = 6^{\frac{1}{3}} e^{i \frac{2 \pi k}{3}} \] what do you get when k= 0 ?

OpenStudy (anonymous):

6^.333 (e^i(0/3) 6^.333 (e^i(0) 6^.333 (e^0) 6^.333 (1) 6^.333 1.817 @phi

OpenStudy (phi):

yes, now what do you get for k=1 ?

OpenStudy (anonymous):

6^.333 (e^i(2pi(1)/3) 6^.333 (e^i(2pi/3) 6^.333 (e^i(2.09) I am not positive what to do with the i(2.09) but on my calculator i did imaginary(2.09) = 0 6^.333 (e^0) 6^.333 (1) 6^.333 1.817 @phi

OpenStudy (phi):

you have to use euler's formula http://en.wikipedia.org/wiki/Euler's_formula because your calculator does not do complex numbers (I don't think) be sure to be in radian mode (or change the radians to degrees) 2 pi/3 in degrees is 2 pi /3 * 180/pi = 120 degrees

OpenStudy (anonymous):

6^.333 (e^i(2pi(1)/3) 6^.333 (e^i(2pi/3) cos^i(2pi/3)=cos(2pi/3)+isin(2pi/3) cos^i(2pi/3)=-.5+i(.866) 6^.333 (e^cos^i(2pi/3)=-.5+i(.866)) am i doing this right? if so what next?

OpenStudy (phi):

6^.333 (e^i(2pi(1)/3) 6^.333 (e^i(2pi/3) you should say (not cos^i(2pi/3)) e^i(2pi/3)=cos(2pi/3)+isin(2pi/3) that means you replace e^i(2pi/3) with the stuff on the right to get 6^.333 ( cos(2pi/3)+isin(2pi/3) ) you should memorize cos(120) is -1/2 and sin(120) is sqrt(3)/2 so you get (remember 6^(1/3) is about 1.817 \[ 1.817\left( - \frac{1}{2} + \frac{\sqrt{3}}{2} i \right) \] or \[ 0.909 ( -1 + \sqrt{3} i) \] or (to be exact) \[ \frac{\sqrt[3]{6}}{2}\left( -1 + \sqrt{3} \ i\right) \]

OpenStudy (phi):

or roughly, -0.909 + 1.574 i

OpenStudy (anonymous):

thanks! and now k=2 1.87 (e^i(2π(2)/3)) 1.87 (e^i(4π/3)) e^i(4π/3)=cos(4π/3+isin(4π/3) e^i(4π/3)= -.5 + -sqrt3/2 i 1.87 (-.5 + -sqrt3/2 i) 0.909(−1-√3i) 3sqrt/2 (−1-√3i) @phi

OpenStudy (phi):

if in the last line 3 sqrt/2 is a typo for \[ \frac{\sqrt[3]{6} }{2} \] then yes, looks good. notice that complex numbers are *complex conjugates* (roots of real numbers always come in complex conjugate pairs)

OpenStudy (anonymous):

yeah thats what i meant and thanks so mucH!